OpenStudy (inowalst):

Jennifer can clean her room in 2 hours, while John can clean the same room in 5 hours. How many hours would it take for them to clean the room together?

OpenStudy (hugsnotughs):

OMG, I learned this. >v> Maybe I can give you an example I did because I learned this a while ago.

OpenStudy (hugsnotughs):

(and because i have to go like in two minutes)

OpenStudy (inowalst):

\[7 \] \[10\] \[\frac{ 7 }{ 10 }\] \[\frac{ 10 }{ 7 }\]

OpenStudy (inowalst):

Those are my options.

OpenStudy (jordanloveangel):

im thinking 7

OpenStudy (inowalst):

What makes you say that?

OpenStudy (jordanloveangel):

http://openstudy.com/study#/updates/53279244e4b087b3d4109a36

OpenStudy (calculusxy):

For this one you need to figure out the LCM , least common mutltiple. There is a multiple that both 2 and 5 overlap on.

OpenStudy (hugsnotughs):

I have to go @inowalst , but here is an example on how you do these kinds of problems: "If Ezra takes 6 hours to build a house and it takes Jack 12 hours to build the same house, how long would it take for them to do it together?" 2. Write a rational equation based upon the word problem you created. 1/6 + 1/12 = 1/x. The numbers are in hours and 1/6 represents Ezra's portion and 1/12 represents Jack's. 3. Solve the rational equation. Show your work. 1/6 + 1/12 = 1/x Find the common demoninator. (1 * 4x)/(6 * 4x) + (1 * 2x)/(12 * 2x) = (1 * 24)/(x * 24) 4x/24x + 2x/2x = 24/24x 4x + 2x = 24 6x = 24 x = 4 Together, it will take Ezra and Jack will take 4 hours to build the house.

OpenStudy (hugsnotughs):

Hope you pass, but I gots to go. ;3

OpenStudy (inowalst):

Alright, thank you!

OpenStudy (hugsnotughs):

Any time. :3

OpenStudy (theraggedydoctor):

dot.